# Causal Predictive Optimization and Generation for Business AI

### v1

@@ 5.1.4. Templates for Recommendations
  • Low Engagement: RTCD = $d$. We recommend reaching out to the client to understand the low engagement in ¡product¿. The current usage is $y$ and we are predicting a drop to $\Delta y$ over the next month.
  • Upsell Flag: RTCD = $d$. We recommend exploring add-on opportunities with this customer as we predict a near-term upsell opportunity worth $\Delta y$.
  • Churn Flag: RTCD = $d$. We recommend connecting with customers to assess churn risks.
@@ 5.2. Template-free explanations with GAI
To enhance the explanation layers and streamline the on-boarding process, we leverage LLM to generate feature explanations for each feature name, and integrate Generative AI (GAI) with sales productivity tools to automate narrative generation and feature categorization. Multiple instance-level explanation algorithms (e.g., CLIME, TE2Rules, Integrated Gradients) are incorporated for calculating the importance of instance-level features (Figure  2). Using an iterative approach, we specify the necessary steps to complete a task and provide a sample output format for GAI to learn. By integrating feature names and explanations along with instance-level feature importance scores into the LLM, we generate detailed instance-level explanations  [^2] For detailed steps, please refer to Appendix A.5.. This includes insights into what changes in each feature might indicate, offering a comprehensive understanding of the data’s impact on model recommendations (Figure  3).
@@
The framework can also integrate multiple components to generate insights and provide conversational experience for user in agentic AI framework (Figure  4). The Content Component processes input data, including tabular, text, and image data, through an Input Data Module. This data is then used in the Insights Module to generate sales funnel-specific, sales call, product and engagement insights, and other relevant insights, which are standardized and normalized. The Agent Component interacts with users, utilizing these insights and generating conversation logs. These logs, including labeled prompt-response pairs, are used in the Fine-Tune module to improve the agentic AI system.
@@ 6. Experiments and Results
We conduct experiments to evaluate our framework using both offline and causal inference. Our evaluation has two main objectives:
  1. (1)
    Evaluate the individual performance of the layers and components using offline data
  2. (2)
    Analyze the business impact of our solution by deploying to a live system and collecting user feedback
6.1. Offline Analysis and Validation
6.1.1. Offline Evaluation Dataset
Training data for our offline evaluation consists of active SME customers, with closed opportunities, during the data collection period. These closed opportunities were either won or disengaged. Our features included customer engagement metrics, historical bookings, product usage, company firmographics, and LinkedIn-only data like hiring trends.
6.1.2. Prediction Layer Implementation
6.1.3. Uplift Model Evaluation
@@ 6.1.5. Ablation analysis for Constrained Optimization
We conducted an ablation analysis of the optimization layer by systematically removing or modifying key components—such as the weighting function, capacity constraints, and recommendation rules. Using six months of backtesting data, we compared the performance of the Full Model (with all components intact) to simplified models across key metrics, including upsell precision, churn precision, churn recovery precision, low engagement precision, total bookings rank, and constraint feasibility (Details in Appendix  A.2).
The Full Model consistently outperformed the ablated models, demonstrating the importance of each component:
@@ 6.1.6. Validation for Contextual Bandit
To evaluate the contextual bandit layer’s performance, we conducted a comprehensive offline analysis. This analysis aimed to validate the bandit’s ability to optimize recommendations based on contextual information and observed rewards. The key areas of analysis include:
  • Feedback Distribution: This analysis examines how frequently each feedback, i.e., action from sales representatives, (DEEP_LINK_CLICKED, NOTIFICATION_DISMISSED, NO_CLICK) was selected during the simulations. The distribution highlights the bandit’s prioritization of high-reward feedback (Figure  7 in Appendix).
  • Feedback Trends Over Time: Temporal trends in Feedback percentages provide insights into the bandit’s progression from exploration to exploitation, i.e., focusing on actions that leads high-reward feedback (Figure  6).
    Figure 6. Feedback Trend Over Time for Thompson Sampling
For the bandit optimization, both Neural Thompson Sampling and Neural UCB approaches were evaluated to compare their effectiveness in terms of policy convergence, cumulative rewards, and their ability to balance exploration and exploitation. The cumulative rewards for both strategies are comparable over multiple iterations (Figure 8 in Appendix). The current system adopts Thompson Sampling due to its low computational overhead per iteration leading to low latency while serving.
6.2. Roll-out and Launch Details
CPOG’s modular and flexible design enables tailored solutions for various customer segments and regional needs. We launched a test within LinkedIn, targeting 25% of sales reps focused on SME customers in North America. Sales reps received personalized recommendations which are delivered in near real-time through CRM applications, seamlessly integrating with sales workflows. Feedback on recommendation relevance is collected to continuously improve future suggestions. Our solution received high user satisfaction among sales reps (details in Appendix  A.3) and delivered significant business impact (+59.96% lift on primary metrics) with minimal disruption to sales workflows and quota achievement.
6.2.1. Challenges to A/B Test
While A/B test is the best practice for impact estimation, several challenges must be considered in our scenario
  1. (1)
    Opportunity Costs: A partial rollout could limit solution availability, resulting in missed opportunities despite its potential to alleviate sales representatives’ pain points with minimal disruption.
  2. (2)
    Business Urgency: Achieving revenue goals amidst macroeconomic pressures necessitates leveraging the solution’s timely insights promptly.
  3. (3)
    Selection Bias: Randomization in treatment group selection may be infeasible, introducing potential selection bias.
  4. (4)
    Quota Equity: Disparities in quotas arising from the test must be mitigated through appropriate quota adjustments during interim sales planning cycles.
  5. (5)
    Statistical Power: Testing with a small group may lack sufficient power to produce statistically significant results.
Given the challenges of conducting a randomized A/B test and the subsequent decision to implement a full rollout after the initial ramp, we implemented an observational causal study to estimate the business impact.
6.2.2. Causal Inference
The treatment and control groups were defined as follows:
@@ 6.2.2. Causal Inference
The units in our treatment group are ¡rep, account¿ pairs because recommendations are generated for only a subset of accounts within each sales reps’ books. We focus exclusively on relevant booking opportunities [^3] Opportunities originating from accounts with valid recommendations, generated within a specified time period, and followed up by sales outreach which makes our treatment period time-dependent. To account for irregular and sparse timing of treatments across multiple periods, we utilized Difference-in-Difference (DiD) (Ashenfelter and Card, 1984) methodology for our causal analysis. DiD estimates the Average Treatment Effect on the Treated (ATT) - mean difference in actual outcomes for the treated units and their counterfactual outcomes.
$ATT=\mathbb{E}[Y(a)-Y(0)|a=1]$
We report the Relative Treatment Effect (RTE) of our ATT estimate - which gives a better intuition of the treatment effect relative to the baseline or control group.
$RTE=ATT/D$
Where $D=\mathbb{E}[Y(a)|a=1]-\mathbb{E}[Y(a)-Y(0)|a=1]$
6.2.3. Assumption Validation
@@ 6.2.4. Experiment Setup
Our online experiments consists of two time periods: (1) pre-intervention - during the second half of 2023 and (2) post-intervention - during the first half of 2024. Our control group consists of candidate accounts outside North America and were matched to the treatment group using Coarsened Exact Matching (CEM) (Iacus et al., 2012). We used a proxy metric, Net Ratio, as the primary response variable for our test period and define Net Ratio as follows:
@@ 7. Conclusion
In this paper, we introduce the Causal Predictive Optimization and Generation (CPOG) framework, a novel approach to sales optimization and business AI. Through offline analysis, online testing and deployment, we demonstrate CPOG’s effectiveness in optimizing sales strategies across multiple metrics. In particular, we highlight the importance of using uplift models to prioritize sales actions, the benefits of constrained optimization in balancing multiple objectives, and the value of providing sales representatives with clear, explainable and actionable recommendations. We give guidelines on how to apply CPOG framework and clear examples of how we implemented it at LinkedIn, along with results and potential measurement approaches in complex applications where A/B testing is not feasible. The framework and the learnings in deployments are broadly applicable to the field, particular B2B and SaaS businesses.
Appendix A Additional Offline Analysis
A.1. Comparison for Uplift Modeling
@@
A.2. Details of Ablation Analysis for Optimization
A.2.1. Objective
@@ A.2.3. Metrics
A.2.4. Results
Table 7 and 8 are results from backtesting data.
@@ A.2.4. Results
A.3. Survey on Explanation Layer
@@
A.4. Explanation tables
Table 7. Feature Name Mapping
Feature Name Expression
producta_l1m Avg. product A usage for last month
productb_l2m Avg. product B usage across last 2 months
Table 8. Basic Information and Thresholds
Feature Name Model Threshold
producta_l1m Treatment Model ¿0
producta_l1m Control Model ¡0
Figure 7. Feedback Distribution for Thompson Sampling
A.5. GAI-Based Instance-Level Explanation Generation Process
A.5.1. Step 1: Feature Grouping using LLM
  • super_name: Meaning of features.
Example Output:
Feature_name super_name ultra_name
MetricA_l1m Metric A in the last month Metric A
MetricA_l2m_l1m Metric A in the last month Metric A
MetricA_l3m Metric A in the last 3 months Metric A
MetricB_l1m Metric B in the last month Metric B
Table 9. Abstracted Feature Grouping Example
A.5.2. Step 2: Combining Feature Importance with Metadata
  1. (1)
  2. (2)
  3. (3)
    Enriching with feature values per account.
Output Schema:
Weight Feature_name Value
0.072 MetricC_l12m 24.0
Table 10. Step 2 output table 1
Super_name Ultra_name customer_urn
Metric C in the last 12 months Metric C urn:li:customer:…
Table 11. Step 2 output table 2
A.5.3. Step 3: Explanation Generation using LLM
Example Output:

### v2

@@ 5.1.2. Templates for Recommendations
  • Low Engagement: RTCD = $d_{1}$. We recommend reaching out to the client to understand the low engagement in ¡product¿. The current usage is $y$ and we are predicting a drop to $\Delta y$ over the next month.
  • Upsell Flag: RTCD = $d_{2}$. We recommend exploring add-on opportunities with this customer as we predict a near-term upsell opportunity worth $\Delta y$.
  • Churn Flag: RTCD = $d_{3}$. We recommend connecting with customers to assess churn risks.
@@ 5.2. Template-free explanations with GAI
To enhance the explanation layers and streamline the on-boarding process, we leverage LLM to generate feature explanations for each feature name, and integrate Generative AI (GAI) with sales productivity tools to automate narrative generation and feature categorization. Multiple instance-level explanation algorithms (e.g., CLIME, TE2Rules, Integrated Gradients) are incorporated for calculating the importance of instance-level features (Figure  2). Using an iterative approach, we specify the necessary steps to complete a task and provide a sample output format for GAI to learn. By integrating feature names and explanations along with instance-level feature importance scores into the LLM, we generate detailed instance-level explanations. This includes insights into what changes in each feature might indicate, offering a comprehensive understanding of the data’s impact on model recommendations (Figure  3). Detais are provided in Appendix A.3.
@@
We also integrate multiple components to generate insights and provide conversational experience for user in agentic AI framework (Figure  4). The Content Component processes input data, including tabular, text, and image data, through an Input Data Module. This data is then used in the Insights Module to generate sales funnel-specific, sales call, product and engagement insights, and other relevant insights, which are standardized and normalized. The Agent Component interacts with users, utilizing these insights and generating conversation logs. These logs, including labeled prompt-response pairs, are used in the Fine-Tune module to improve the agentic AI system.
@@ 6. Experiments and Results
6.1. Offline Analysis
We evaluate the performance of each layers of CPOG in offline Analysis. We provide high level summary here. Details can be found in Appendix B.
6.1.1. Offline Evaluation for Prediction Models
We evaluated 4 uplift estimators: S-learner, T-learner, X-learner, and DR-Learner, by ranking customers into uplift deciles(Gutierrez and Gérardy, 2017) and examining uplift curves (Figure 5), while also considering prediction stability and score interpretability. Although the S- and X-learners produced sensible ITE distributions, we ultimately selected the T-learner for its consistently stable decile bins and ITE estimates across multiple runs.
For customer engagement, we modeled monthly product adoption ($pa$) and utilization ($pu$) and measured performance via Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE); the results for models $f_{pu}$ and $f_{pa}$ are summarized in Table 3.
@@ 6.1.2. Ablation analysis for Constrained Optimization
We conducted an ablation analysis of the optimization layer by systematically removing or modifying key components—such as the weighting function, capacity constraints, and recommendation rules. Using six months of backtesting data, we compared the performance of the Full Model (with all components intact) to simplified models across key metrics, including upsell precision, churn precision, churn recovery precision, low engagement precision, total bookings rank, and constraint feasibility. The Full Model consistently outperformed the ablated models, demonstrating the importance of each component. (Details in Appendix  B.2)
@@ 6.1.3. Validation for Contextual Bandit
We validated the contextual bandit layer via offline simulations that assessed its recommendation quality and learning behavior. First, we analyzed the feedback distribution which illustrates the bandit’s ability to prioritize recommendations that results in high-reward signals. We then examined feedback trends over time, revealing the transition from exploration to exploitation as the model increasingly favors actions yielding greater rewards. Finally, we compared cumulative rewards for Neural Thompson Sampling and Neural UCB to evaluate policy convergence and exploration–exploitation balance; both methods achieved similar overall returns, but we selected Thompson Sampling for production due to its lower per-iteration computational cost and resulting low serving latency. (Details in Appendix B.3)
6.2. Online Experiment
We launched an online test within LinkedIn, targeting sales reps that focused on SME customers in North America. Sales reps received personalized recommendations which are delivered in near real-time through CRM applications. Our solution received high user satisfaction among sales reps (details in Appendix  C.2) and delivered significant business impact (+59.96% lift on primary metrics) with minimal disruption to sales workflows and quota achievement.
6.2.1. Challenges on randomized A/B Test
While randomized A/B test is the best practice for impact estimation, several challenges must be considered in our scenario: (i) Opportunity Costs, A partial rollout could limit solution availability, resulting in missed opportunities despite its potential to alleviate sales representatives’ pain points with minimal disruption; (ii) Business Urgency, Achieving revenue goals amidst macroeconomic pressures necessitates leveraging the solution’s timely insights promptly; (iii) Selection Bias: Randomization in treatment group selection may be infeasible, introducing potential selection bias; (iv) Quota Equity, Disparities in quotas arising from the test must be mitigated through appropriate quota adjustments during interim sales planning cycles; (vi) Statistical Power, Testing with a small group may lack sufficient power to produce statistically significant results.
6.2.2. Observational Study
Given the challenges of conducting a randomized A/B test, we implemented an observational causal study to estimate the business impact. The treatment and control groups were defined as follows:
@@ 6.2.2. Observational Study
The units in our treatment group are ¡rep, account¿ pairs because recommendations are generated for only a subset of accounts within each sales reps’ books. We focus exclusively on relevant booking opportunities [^2] Opportunities originating from accounts with valid recommendations, generated within a specified time period, and followed up by sales outreach which makes our treatment period time-dependent. To account for irregular and sparse timing of treatments across multiple periods, we utilized Difference-in-Difference (DiD) (Ashenfelter and Card, 1984) methodology for our causal analysis. Formally, the DiD estimator is defined as,
$\widehat{\tau}_{\text{DiD}}=\bigl{(}\bar{Y}_{\text{treat},1}-\bar{Y}_{\text{treat},0}\bigr{)}-\bigl{(}\bar{Y}_{\text{ctrl},1}-\bar{Y}_{\text{ctrl},0}\bigr{)}$
which measures the treatment effect under the parallel‐trends assumption. We can calculate the relative treatment effect by,
$\text{RTE}=\frac{\widehat{\tau}_{\text{DiD}}}{\bar{Y}_{\text{ctrl},1}-\bar{Y}_{\text{ctrl},0}},$
Details on assumption validation are in Appendix C.1.
@@ 6.2.3. Experiment Setup
Our online experiments consists of two time periods: (1) A/A - 6 months before intervention and (2) observational (non-randomized) A/B - 6 months after experiment started. Our control group consists of candidate accounts outside North America and were matched to the treatment group using Coarsened Exact Matching (CEM) (Iacus et al., 2012). We used a proxy metric, Net Ratio, as the primary response variable for our test period and define Net Ratio as follows:
@@ 7. Conclusion
In this paper, we introduce the Causal Predictive Optimization and Generation (CPOG) framework, a novel approach to sales optimization and business AI. Through offline analysis, online testing and deployment, we demonstrate CPOG’s effectiveness in optimizing sales strategies across multiple metrics. In particular, we highlight the importance of using uplift models to prioritize sales actions, the benefits of constrained optimization in balancing multiple objectives, and the value of providing sales representatives with clear, explainable and actionable recommendations. We give guidelines on how to apply CPOG framework and clear examples of how we implemented it at LinkedIn, along with results and potential measurement approaches in complex applications where randomized A/B testing is not feasible. The framework and the learnings in deployments are broadly applicable to the field, particular B2B and SaaS businesses.
Appendix A Additional Details on Method
A.1. Uplift Modeling
@@
A.2. Explanation tables
Table 5. Feature Name Mapping
Feature Name Expression
producta_l1m Avg. product A usage for last month
productb_l2m Avg. product B usage across last 2 months
Table 6. Basic Information and Thresholds
Feature Name Model Threshold
producta_l1m Treatment Model ¿0
producta_l1m Control Model ¡0
A.3. GAI-Based Instance-Level Explanation Generation Process
A.3.1. Step 1: Feature Grouping using LLM
  • super_name: Meaning of features.
Example Output:
Feature_name super_name ultra_name
MetricA_l1m Metric A in the last month Metric A
MetricA_l2m_l1m Metric A in the last month Metric A
MetricA_l3m Metric A in the last 3 months Metric A
MetricB_l1m Metric B in the last month Metric B
Table 7. Abstracted Feature Grouping Example
A.3.2. Step 2: Combining Feature Importance with Metadata
  1. (1)
  2. (2)
  3. (3)
    Enriching with feature values per account.
Output Schema:
Weight Feature_name Value
0.072 MetricC_l12m 24.0
Table 8. Step 2 output table 1
Super_name Ultra_name customer_urn
Metric C in the last 12 months Metric C urn:li:customer:…
Table 9. Step 2 output table 2
A.3.3. Step 3: Explanation Generation using LLM
Example Output:
Appendix B Additional Offline Analysis
B.1. Details of Offline Evaluation for Prediction
B.1.1. Dataset and Features
Training data for offline evaluation consists of active SME customers, with closed opportunities, during the data collection period. These closed opportunities were either won or disengaged. Our features included customer engagement metrics, historical bookings, product usage, company firmographics, and LinkedIn-only data like hiring trends.
B.1.2. Uplift Model Evaluation
B.1.3. Customer Engagement Models Evaluation
B.2. Details of Ablation Analysis for Optimization
B.2.1. Objective
@@ B.2.3. Metrics
B.2.4. Results
Table 10 and 11 are results from backtesting data.
@@ B.2.4. Results
B.3. Details of Validation Analysis for Contextual Bandit
To evaluate the contextual bandit layer’s performance, we conducted offline analysis. This analysis aimed to validate the bandit’s ability to optimize recommendations based on contextual information and observed rewards. The key areas of analysis include:
  • Feedback Distribution: This analysis examines how frequently each feedback, i.e., action from sales representatives, (DEEP_LINK_CLICKED, NOTIFICATION_DISMISSED, NO_CLICK) was selected during the simulations. The distribution highlights the bandit’s prioritization of high-reward feedback (Figure  6).
  • Feedback Trends Over Time: Temporal trends in Feedback percentages provide insights into the bandit’s progression from exploration to exploitation, i.e., focusing on actions that leads high-reward feedback (Figure  7).
Figure 6. Feedback Distribution for Thompson Sampling
Figure 7. Feedback Trend Over Time for Thompson Sampling
For the bandit optimization, both Neural Thompson Sampling and Neural UCB approaches were evaluated to compare their effectiveness in terms of policy convergence, cumulative rewards, and their ability to balance exploration and exploitation. The cumulative rewards for both strategies are comparable over multiple iterations (Figure 8). The current system adopts Thompson Sampling due to its low computational overhead per iteration leading to low latency while serving.
Appendix C Additional Details for Online Experiment
C.1. Causal Inference Analysis and Assumption Validation
C.2. Survey on Explanation Layer